OpenStudy (anonymous):

What is the sum of the arithmetic sequence 6, 14, 22, if there are 26 terms

OpenStudy (amistre64):

do you know the formula to find any term in the sequence?

OpenStudy (anonymous):

isnt it like an=a1(r)^2-1?

OpenStudy (amistre64):

it "like" that yes, but that ones looks more like the GEOmetric sequence formula the one for arithmetic is: an = a1 + d(n-1)

OpenStudy (anonymous):

ok so what do I do?

OpenStudy (amistre64):

you have a1, you know how to find any term now, and you want the 26th term. All your missing is the common difference "d" between each term. but, your given 3 terms to begin with, so number do you have to add to the firs term to get the second term? a1 + d = a2

OpenStudy (anonymous):

the difference is 8 so what do i do now?

OpenStudy (amistre64):

we have all the pieces we need to find the 26th term, then we can find the sum

OpenStudy (amistre64):

a26 = 6 + 8(26-1)

OpenStudy (anonymous):

a26=206?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

how we have all the parts we need for the next formula n(a1 + a26)/2 ; where n = 26 and we know a1 and a26 26(6+208)/2

OpenStudy (amistre64):

206 .... not 208 :)

OpenStudy (amistre64):

26(6+206)/2 = 26(3+103) = 26(106)

OpenStudy (anonymous):

so it 2782?

OpenStudy (anonymous):

wait no what is it?

OpenStudy (anonymous):

ooooo 2756

OpenStudy (amistre64):

i get 2756 with the correct values :)

OpenStudy (amistre64):

it simpler when you know why the formulas work, they realy are quite intuitive. but if all you see is a formula and have to memorize the formula, it can be challenging

OpenStudy (anonymous):

THANKS:)

OpenStudy (amistre64):

youre welcome